OpenStudy (anonymous):

Solve 4/5 x- 3 3/4 x= 2 1/2?

OpenStudy (anonymous):

\[\frac{ 4x }{ 5}-\frac{ 15x }{ 4 } =\frac{ 5 }{2}\] so this is simpllified eqn

OpenStudy (anonymous):

if this is correct then i will help u solve this

OpenStudy (anonymous):

It seems to be correct, I mean it would make sense

OpenStudy (anonymous):

hmm then take the lcm of LHS \[\frac{(4)(4)(x)}{ 20 }- \frac{ 15(5)(x) }{ 20 }=\frac{ 5 }{ }\]

OpenStudy (anonymous):

\[\frac{ 16x-75x }{ 20 }= 5/2\]

OpenStudy (anonymous):

 cross multiply